OpenStudy (anonymous):

I just need someone to check my work!x-y is less than or equal to 7 x+2y <5 @texaschic101

OpenStudy (texaschic101):

what are you trying to do here ?

OpenStudy (anonymous):

Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system

OpenStudy (anonymous):

6,1 8,-1 6,2 6,-2

OpenStudy (anonymous):

I think its 6,-2

OpenStudy (anonymous):

What do you think it is? @texaschic101

OpenStudy (texaschic101):

Is the problem written correctly ? ...

OpenStudy (anonymous):

No omg I meant to put that x-y is greater than or equal to

OpenStudy (texaschic101):

x - y >= 7 6 - (-2) >= 7 6 + 2 >= 7 8 >= 7 (correct) x + 2y < 5 6 + 2(-2) < 5 6 - 4 < 5 2 < 5 (correct) your right...answer is (6,-2)

OpenStudy (anonymous):

CAn you check 2 more please???

OpenStudy (texaschic101):

I am sorry but I have to go somewhere right now....I will probably be back on later...sorry :(

OpenStudy (anonymous):

Its ok you were a huge help thanks!!!

OpenStudy (texaschic101):

YW :)